2/3x + 3/5= 1/5 - 2x
Answer by stanbon(75887)

2/3x + 3/5= 1/5 - 2x
--------------
Multiply thru by 15 to get::
10x + 9 = 3 - 30x
--------
40x = -6
x = -3/20
------------------
Cheers,
Stan H.
--------------